If the sum of the interior angles of a polygon is 12 right angles, find the number of sides it has.
Answer by ikleyn(52778)

.
(n-2)*180° = 12*90°,

(n-2)*180° = 6*180°,

n-2 = 6,

n = 8.

All that you need to know is that the sum of interior angles of an n-sided polygon is (n-2)*180°.